Two equal forces(F) act at a point inclined to each other at an angle of 120* what is the resultant of their magnitude
 

Grade:11

2 Answers

SWAMY
4 years ago
Hmm, you want fastest method . 
you can do it fast when you know some special result. 
There are ..... 
1. When two equal vectors A are inclined 90° then, resultant = √2A 
2. When angle between two equal vectors A is 120° , then resultant = A 
3. When angle between two equal vectors A is 180° , then resultant = 0 
4. When angle between two equal vectors A is 60° , then resultant = √3A 
Now, come to the question, 
Your question is " two equal force inclined 120° , 
So, resultant = P 
Just 1 second is enough for answering. 
Well you know method is Resultant = √{P² + P² + 2P.PCos120°} = P
 
 
Saurabh Koranglekar
4 years ago
Dear student

Let F1 = F i
so F2 will be = F(cos 120)i + F sin(120)


R = (F +(-0.5)F)i + F(0.866)j
R = 0.5 F i + 0.866F j
|R|≈ F

Regards